### v2

@@ 2.7 Effect of Chunked Contexts
2.8 Ablations of EHR Serialization
To assess the impact of different components of the EHR serialization, we conducted a series of ablation studies for both LLM-embedding models, as summarized in Table 6. The removal of instructions had a notable effect on performance for both GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, leading to a decrease from 0.774 (0.749 - 0.798) to 0.725 (0.700 - 0.750) and from 0.742 (0.714 - 0.769) to 0.727 (0.700 - 0.755) in averaged AUROC, respectively. The most significant drop was observed in tasks related to operational outcomes and lab result prediction, suggesting that the instructions played a critical role in guiding the model to focus on relevant clinical information. The inclusion of aggregated information, comprising four body metrics, six vital signs, and 14 lab values, also had a major impact on predictive performance. Removing this information substantially reduced the accuracy of lab test predictions, with performance dropping from 0.867 (0.860 - 0.874) to 0.713 (0.701 - 0.726) for GTE-Qwen2-7B and from 0.777 (0.766 - 0.787) to 0.705 (0.692 - 0.718) for LLM2Vec-Llama-3.1-8B. This highlights the importance of the most recent lab values in predicting future lab results. The effect was particularly pronounced for GTE-Qwen2-7B in predicting conditions such as thrombocytopenia, hyperkalemia, and hyponatremia. However, for hypoglycemia and anemia, GTE-Qwen2-7B performed on par with CLIMBR-T-Base, even though the lab values were still part of the serialized representation (see Fig. 7 and Fig. 8 for task specific performance). Interestingly, the removal of aggregated information led to slight performance improvements in predicting operational outcomes and the assignment of new diagnoses, suggesting that in some cases, a more focused representation may be beneficial. Visit-related information proved to be particularly critical for operational outcome prediction, as removing it led to a performance drop from 0.844 (0.821 - 0.867) to 0.752 (0.728 - 0.776) for GTE-Qwen2-7B and from 0.787 (0.762 - 0.812) to 0.734 (0.709 - 0.759) for LLM2Vec-Llama-3.1-8B. Notably, the removal of visit information resulted in a substantial performance increase for LLM2Vec-Llama-3.1-8B predicting lab values, from 0.777 (0.766 - 0.787) to 0.862 (0.855 - 0.869). This suggests that the model struggles to extract relevant information when presented with a large volume of contextual data, a challenge that GTE-Qwen2-7B handles more effectively. On the other hand, the removal of demographic information, conditions, medications, and procedures had only a minor effect on performance, indicating that these features were either redundant or less relevant for the prediction tasks. While the influence of aggregated information and visit details was clear for predicting operational outcomes and lab results, the ablation studies showed less impact on tasks related to new diagnoses and chest X-ray prediction. This suggests that these latter task groups rely more on a broader set of contextual factors rather than any single component of the serialization.
@@ 3 Discussion
Our study demonstrates that general-purpose LLM-embedding models, originally pre-trained on extensive natural language corpora, can be repurposed as robust foundation models for EHR prediction tasks. Specifically, models like GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B not only surpassed a strong counts-based baseline but, in several clinical domains, matched or even exceeded the performance of a dedicated EHR foundation model (CLIMBR-T-Base) [13, 30]. Notably, this is despite CLIMBR-T-Base being trained on data from the same hospital system as the EHRSHOT database, underscoring the strong generalization capabilities of LLM-based embeddings. This is especially significant given the inherent challenges of limited and heterogeneous EHR data. Across 15 diverse clinical tasks, including few-shot scenarios, our findings indicate that the transferable knowledge acquired during large-scale text pretraining enables these models to effectively capture and encode complex clinical information. Furthermore, our analysis shows that performance scales with both model size and architectural improvements, suggesting that future advancements in general-purpose LLMs may further enhance their applicability in healthcare settings. In contrast, traditional EHR-specific models face scaling limitations due to restricted data availability and the necessity for specialized architectures [34]. Collectively, these results highlight that repurposing LLMs as embedding generators offers a powerful, flexible, and scalable alternative for clinical prediction tasks.
Our experiments suggest that several key factors contribute to the success of LLM-embedding models for encoding EHR data. First, providing clear, task-specific instructions within the serialized EHR text was crucial, as it guided the models to focus on the most clinically relevant sections, particularly benefiting tasks that require attention to specific input segments. This design leverages the general-purpose pretraining and further instruction-based fine-tuning of LLM-embedding models [26, 29], enabling them to identify and extract meaningful patterns from heterogeneous and complex EHR inputs. Second, the inclusion of aggregated clinical data, especially semantic LOINC-coded events and detailed visit-related information, was strongly associated with improved performance on operational outcomes and lab test predictions. In contrast, other components of the EHR did not show as clear a relationship with predictive accuracy. The effective handling of context size also plays a significant role. While models like GTE-Qwen2-7B can fully utilize contexts up to 4,096 tokens, our findings show that performance degrades with overly short or excessively long inputs. Short contexts risk omitting recent, crucial clinical encounters, whereas overly long contexts can dilute important token-level details. In cases where models face challenges with longer inputs such as LLM2Vec-Llama-3.1-8B our chunking experiments indicate that segmenting the input and averaging the resulting embeddings can mitigate this issue. Finally, combining LLM embeddings with those from an EHR-specific foundation model (CLMBR-T-Base) led to a considerable boost in performance. This improvement suggests that LLMs capture orthogonal and complementary clinical information. One explanation for this effect could be the limited context size of CLMBR, restricted to 496 medical codes, which prevents it from fully leveraging all information present in our EHR serialization [13]. Alternatively, the boost may stem from the general-purpose pretraining on vast text corpora, which provides LLMs with domain-agnostic knowledge that remains untapped in models trained exclusively on EHR data.
@@
Unlike count-based models and specialized EHR foundation models, LLM-based models are agnostic towards specific coding systems and data formats. Traditional EHR models depend on predefined vocabularies such as SNOMED, LOINC, and ICD, which restricts their applicability across diverse healthcare settings. In contrast, LLMs operate on raw textual representations, enabling them to interpret any code that can be mapped to a human-readable form. This flexibility is particularly valuable given the persistent challenges in EHR data interoperability, stemming from variable coding practices, privacy constraints, and regulatory limitations, which often hinder the aggregation of large, standardized datasets for model pretraining [35]. Moreover, since LLMs are pre-trained on broad and diverse text corpora, including medical literature and case reports, they are well-equipped to capture the semantics of rare or underrepresented clinical concepts [36]. This capacity mitigates the limitations of count-based models that typically filter out infrequent events, ensuring that even rare but clinically significant phenomena are effectively encoded [37]. Additionally, the ability of LLMs to process both structured and unstructured data paves the way for integrated, multimodal representations that could further enhance clinical decision support systems [38].
@@
Our findings highlight that LLM-embedding methods merge the strengths of count-based models, specialized EHR models, and the expansive generalization capabilities of large language models. While count-based and specialized EHR foundation models can be tuned to output well-calibrated probabilities, thereby enhancing trust in clinical applications, they are inherently limited by their reliance on predefined vocabularies and constrained training data. In contrast, LLMs can process any textual representation, enabling them to effectively handle rare or unseen codes, although they typically produce unstructured outputs that challenge clinical grounding and calibration [23]. LLM-embedding methods bridge this gap by leveraging the representational power of LLMs while maintaining compatibility with traditional predictive modeling frameworks. By transforming EHR data into structured text and embedding it using general-purpose LLMs, we enable the application of logistic regression or other conventional machine learning classifiers to make clinically meaningful predictions. This approach preserves the calibration advantages of traditional models while benefiting from the rich contextual understanding of LLMs.
3.1 Limitations
@@
This study has several limitations. First, our approach relies on a subjectively designed EHR serialization that we deemed to capture the most medically relevant information. This design choice may introduce bias when comparing LLM-based embedding models with dedicated EHR foundation models that operate on raw data. In addition, the performance of the LLM-embedding models is sensitive to the specific content and instructions provided in the serialized text, which may limit reproducibility and generalization ability. Although these models achieve competitive predictive performance across diverse clinical tasks, including few-shot scenarios, their substantially larger parameter counts lead to longer computation times and higher resource usage. Moreover, by relying on LLM embedding methods, we must train a downstream classifier from scratch, thereby forgoing the inherent zero-shot or few-shot prompting capabilities of LLMs. Finally, our serialization was limited to 4,096 tokens to manage runtime constraints, potentially omitting valuable long-range historical information, and our evaluation on a single institutional dataset may limit broader applicability across diverse healthcare systems.
3.2 Future work
@@
Future research should explore serialization‐free approaches that allow LLMs to process raw EHR data directly, thereby reducing potential biases introduced by manual text transformation. Integrating zero-shot and few-shot prompting into the LLM-based embedding framework could further enhance model flexibility and reduce dependency on downstream training. It will also be important to develop strategies to extend the effective context window beyond 4.096 tokens to capture more comprehensive patient histories. Moreover, investigating techniques for distilling large LLMs into smaller, more efficient models may enhance their practical applicability in clinical settings. Finally, expanding evaluations to multi-institutional datasets and examining how complementary insights from both domain-specific EHR models and general-purpose LLMs can be synergistically combined will be critical for advancing the development of robust, scalable EHR foundation models.
@@ 4 Methods
4.1 EHR Database and Prediction Task
The EHR data utilized in our experiments is from the EHRSHOT benchmark for few-shot evaluation of EHR foundation models [30]. We obtained version 2.1 of the dataset, which is accessible via gated access under a research data use agreement. This dataset comprises longitudinal records for 6,739 patients, 921,499 visits, and 41,661,637 clinical events collected between 1990 and February 8th, 2023. Each clinical event is linked to a specific patient and includes information such as start time, end time, a semantic code, a value, unit, visit ID, and the corresponding OMOP source table. We used the official ehrshot-benchmark repository111Github repository: https://github.com/som-shahlab/ehrshot-benchmark as a starting point to design our experiments enabling to us to build on existing functionalities and to facilitate comparisons with prior methods. The benchmark uses the Framework for Electronic Medical Records (FEMR) 222Github repository: https://github.com/som-shahlab/femr, which provides Python classes for efficient loading and processing of EHR data. All extensions and experiments conducted for this paper have been made publicly available via our GitHub repository: https://github.com/stefanhgm/ehrshot-benchmark. The EHRSHOT benchmark defines a rigorous evaluation including 15 clinical predictions tasks categorized into four groups: operational outcomes, anticipating lab test results, assignment of new diagnoses, and anticipating chest X-ray findings [30]. Task labels are derived from clinical events, resulting in significant variations in task-specific sample sizes. For instance, frequent events like lab tests provide considerably more examples compared to rarer events such as new diagnoses. The benchmark focuses on analyzing model performance in a few-shot setting, which is particularly relevant for large pre-trained foundation models [14] due to their ability to generalize from limited training data. To this end, the benchmark defines evaluation settings with a constrained number of training and validation examples. Specifically, for k in 1, 2, 4, 8, 12, 16, 24, 32, 48, 64, 128, the benchmark uses k positive and k negative training examples, along with k positive and k negative validation examples, to train and tune supervised classifiers. Testing is always performed on the full set of examples. The classifiers evaluated within the EHRSHOT framework include logistic regression, random forests, and gradient boosting machines [39]. Performance is reported using the area under the receiver operating characteristic curve (AUROC) and the area under the precision-recall curve (AUPRC). For few-shot settings, we average the results over five runs with different seeds and compute bootstrapped 95% confidence intervals [30]. Macro averages are reported for each task group, and an overall macro average is provided across all groups.
@@ 4.2 EHR Text Serialization
To leverage LLM-embedding models for representing EHR records, we serialized the records into textual formats. We had to limit the length of the serializations to 4.096 tokens (approximately 16.000 characters) to carry out all experiments on the available computing infrastructure (see Section 4.6). Two experimental runs were conducted with serializations extending up to 8.192 tokens. The primary goal was to create a detailed and informative serialization requiring minimal preprocessing while ensuring that medically relevant information appeared early in the text. This approach mitigated truncation risks in lengthy records, preserving critical details even when older entries were omitted.
@@
To convert the visits and clinical events in the EHRSHOT dataset into text, we leveraged the semantic information embedded in the dataset. Each clinical event was labeled using the format “ontology/code”. EHRSHOT provided a set of prepared ontologies for resolving concept codes into their descriptions, which we incorporated. An analysis was performed to evaluate the utilization of these ontologies for all events of a subset of 200 patients across the task groups for operational outcomes and new diagnoses, covering 2.968 labels. We identified the following ontologies: Logical Observation Identifiers Names and Codes, SNOMED, RxNorm, CPT4, Domain, CARE_SITE, RxNorm Extension, Medicare Specialty, ICD10PCS, CMS Place of Service, Cancer Modifier, ICD9Proc, CVX, ICDO3, HCPCS, OMOP Extension, Condition Type. We excluded ontologies containing only a single value (Domain, Medicare Specialty, CMS Place of Service, OMOP Extension, and Condition Type). Codes of the ontologies CPT4, CARE_SITE, ICD10PCS, Cancer Modifier, CVX, and ICDO3 were not resolved with the provided ontology. Cancer Modifier code contained UICC cancer stages that we parsed manually. For CPT4, ICD10PCS, and CVX we used custom mapping files that we manually added.333We downloaded CPT4 from https://gist.github.com/lieldulev/439793dc3c5a6613b661c33d71fdd185, ICD10PCS from https://hcup-us.ahrq.gov/toolssoftware/procedureicd10/procedure_icd10_archive.jsp, and CVX from https://www2a.cdc.gov/vaccines/iis/iisstandards/vaccines.asp?rpt=cvx. We excluded CARE_SITE and ICDO3 since we found no way to resolve these to useful description.
@@
Various approaches exist for serializing structured data, including row-by-row serialization [40], template-based methods [41], or structured data formats like JSON, HTML, and Markdown [42]. We used Markdown due to its minimal overhead and overall benefits of using a structured data input format for LLMs [43]. To harmonize dates, all timestamps were normalized relative to January 1, 2024, designated as the prediction reference time. Serialization began with patient demographics, typically the first event for each patient, where birthdates were converted into ages (in years) for simplicity. Since 65% of the dataset comprised time-series data encoded via LOINC, which we found imbalanced, we aggregated LOINC-coded events. Using the same patient subset as the ontology analysis, we identified the most frequent codes and categorized them into vital signs, body metrics, and lab values, selecting 24 key medical concepts. To avoid duplicates, we merged synonymous codes (see Table 7). The last three values of each concept are given, and we filtered implausible values. To further enrich the text representation, we manually added default units and assessments (low, normal, high) based on standard ranges (see Table 7). Following the aggregated data, a summary of all visits was included to address the potential truncation of older visits. Events not associated with visits were then presented, using the same aggregation logic to display the last three values where applicable. Finally, a detailed chronological presentation of all visits was included, with events categorized into conditions (SNOMED, Visit, Cancer Modifier, CVX, HCPCS), medications (RxNorm, RxNorm Extension), and procedures (CPT4, ICD10PCS, ICD9Proc).

### v3

@@ 2.3 LLM Embeddings Achieve Strong Performance in Low-Data Regimes
2.4 Effect of Serialization Components on LLM-Based EHR Embedding Performance
To evaluate the contribution of individual serialization components, we conducted ablation studies for both LLM embedding models (see Table 5). Replacing task-specific instructions with a generic prompt led to a marked performance decline, reducing average AUROC from 0.755 (0.730-0.780) to 0.739 (0.713-0.765) for GTE-Qwen2-7B and from 0.734 (0.707-0.762) to 0.724 (0.696-0.752) for LLM2Vec-Llama-3.1-8B. Removing the instruction entirely caused a further drop for GTE-Qwen2-7B to 0.724 (0.698-0.749), whereas LLM2Vec-Llama-3.1-8B showed minimal change, suggesting greater reliance on instruction prompting by GTE-Qwen2-7B. The most significant drop was observed in lab result prediction tasks, highlighting the importance of guidance via task-specific instructions. Aggregated measurements (body metrics, vital signs, and lab values) were crucial, especially for lab test prediction, where their removal reduced AUROC from 0.865 (0.858-0.873) to 0.714 (0.701-0.726) for GTE-Qwen2-7B and from 0.778 (0.768-0.789) to 0.706 (0.693-0.719) for LLM2Vec-Llama-3.1-8B. This underscores the predictive value of recent lab results and their effective encoding by the models. For conditions like thrombocytopenia, hyperkalemia, and hyponatremia, the performance drop was especially pronounced without this information for GTE-Qwen2-7B (see Fig. 8). Interestingly, removing aggregated values slightly improved new diagnosis prediction for GTE-Qwen2-7B, suggesting a more focused input may benefit some tasks. Visit-level information had a strong influence on operational outcome tasks. Its removal led to AUROC reductions from 0.771 (0.747-0.795) to 0.752 (0.728-0.776) for GTE-Qwen2-7B and from 0.753 (0.727-0.778) to 0.734 (0.709-0.759) for LLM2Vec-Llama-3.1-8B. However, for LLM2Vec-Llama-3.1-8B, removing visit data unexpectedly improved lab result prediction, suggesting possible difficulty in filtering relevant signals from verbose input. Removing demographics, conditions, medications, and procedures had only minor effects, indicating these features were either redundant or contributed little predictive signal in isolation. While aggregated information and visit data were key drivers for lab and operational outcomes, prediction of new diagnoses and chest X-ray findings was less sensitive to individual serialization components. This suggests these tasks rely on broader contextual patterns rather than specific structured elements.
2.5 Effect of Temporal Scope and Context Length on Embedding Effectiveness
To assess model sensitivity to input length and recency, we performed ablations on temporal window, context size, and chunking using EHRSHOT data. Both LLM embedding models showed improved performance when restricted to more recent data, with the 1-month window outperforming longer histories. This suggests that recent clinical information may carry the most predictive signal, while older records may contribute less relevant or potentially distracting details (see Fig. 6 and Section A.2). The models differed in optimal context size: GTE-Qwen2-7B achieved the best performance at the maximum 4,096-token context, whereas LLM2Vec-Llama-3.1-8B performed best at 2,048 tokens (Fig. 6 and Table 10). Both models suffered large drops at 512 tokens, a common constraint in older language models [37]. To test whether models processed the entire input in context, we evaluated chunked context representations, where fixed-size segments (e.g., 512–2,048 tokens) of a 4,096-token input were embedded separately and averaged (see Table 11). GTE-Qwen2-7B showed little degradation from chunking, indicating that simple averaging can be sufficient to preserve performance even under token limits. In contrast, LLM2Vec-Llama-3.1-8B benefited from chunking, suggesting it struggles with longer contiguous sequences and prefers shorter inputs. These results highlight differences in model-specific sequence processing and suggest practical approaches for adapting to context length constraints. Full tables and additional details are available in Section A.3 and Section A.4.
@@ 3 Discussion
Our study shows that general-purpose LLM embedding models, pre-trained on large-scale natural language corpora, can be effectively repurposed for clinical prediction using EHR data. Models such as GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B outperformed a strong count-based baseline and, in several clinical domains, matched or surpassed the dedicated EHR foundation model CLMBR-T-Base [13, 31]. This is notable, as CLMBR-T-Base was trained on data from the same health system as EHRSHOT. The strong performance of the LLM-based encoders on the UKB dataset highlights their generalization capabilities. GTE-Qwen2-7B consistently outperformed the domain-specific foundation model with at least 16 training examples, while LLM2Vec-Llama-3.1-8B achieved similar results. These results are particularly notable given the challenges posed by EHR data, including limited size, heterogeneity, and domain shift. Across a broad set of tasks, including few-shot settings, our findings indicate that knowledge acquired during large-scale text pre-training enables LLMs to extract clinically meaningful patterns from serialized EHRs. Model performance improved with larger model sizes and more advanced architectures, suggesting further gains are likely with future LLM development. In contrast, domain-specific EHR models remain constrained by the scarcity and fragmentation of large, diverse clinical datasets. Overall, our findings support general-purpose LLMs as scalable, adaptable, and high-performing alternatives for EHR representation and clinical prediction.
Our experiments suggest that several key factors contribute to the success of LLM embedding models for encoding EHR data. First, providing clear, task-specific instructions in the serialized EHR text was crucial, guiding models to focus on clinically relevant sections and improving tasks requiring targeted attention. This takes advantage of the instruction-tuned nature of LLM embedding models [27, 30], enabling them to extract meaningful patterns from heterogeneous EHR inputs. Second, aggregating structured data, particularly recent lab results and visit-level information, substantially improved performance, especially for lab test and operational outcome prediction. In contrast, components like demographics, medications, and condition codes contributed less consistently to model accuracy. Context size was also critical: GTE-Qwen2-7B effectively handled up to 4,096 tokens, while both models performed poorly with very short or overly long inputs. Intermediate time windows, especially those covering one month prior to prediction, yielded optimal performance, suggesting that recent clinical data carries more predictive value than long-term history for many tasks. For instance, lab values reflect dynamic physiological states and may lose relevance outside of a narrow temporal window. This indicates that LLM embeddings are particularly effective at encoding temporally focused, task-relevant clinical signals. Notably, GTE-Qwen2-7B demonstrated greater robustness to extended inputs, highlighting its stronger capacity to extract salient information from longer contexts. Finally, combining LLM embeddings with CLMBR-T-Base representations improved overall performance, suggesting that these models capture complementary dimensions of patient history. This synergy may stem from CLMBR-T-Base’s limited code-based input window of 496 medical codes [13], or from the broader world knowledge encoded in LLMs that is absent in domain-specific models.
@@
Unlike count-based models and specialized EHR foundation models, LLM-based methods are agnostic to coding systems and data formats. Traditional EHR models rely on predefined vocabularies such as SNOMED, LOINC, and ICD, which can limit their applicability across healthcare settings. For instance, CLMBR-T-Base processes only 65,536 codes (26,249 unique), selected based on their entropy from the EHRSHOT dataset [31]. This restriction may lead to the exclusion of clinically important information when applied to external datasets with unseen codes. In contrast, LLMs operate on textual inputs, allowing them to interpret any clinical code mapped to a human-readable description. This flexibility is particularly valuable given the ongoing challenges in EHR interoperability due to varying coding practices, privacy constraints, and regulatory barriers, which often hinder the aggregation of large, standardized datasets for model pre-training [38]. Moreover, since LLMs are pre-trained on broad and diverse text corpora, including medical literature and case reports, they are well-equipped to capture the semantics of rare or underrepresented clinical concepts [34]. This mitigates the limitations of count-based models that often exclude infrequent events, ensuring that even rare but clinically significant phenomena are effectively encoded [39]. Finally, the ability to represent both structured and unstructured information makes LLMs well-suited for future multimodal EHR applications, including those that integrate clinical notes, diagnostic reports, and imaging metadata [40].
@@
Our findings highlight that LLM embedding methods combine the strengths of count-based models, specialized EHR models, and the generalization capabilities of LLMs. While count-based and specialized EHR foundation models can be tuned to output well-calibrated probabilities, thereby enhancing trust in clinical applications, they are limited by predefined vocabularies and constrained training data. In contrast, LLMs can process any textual representation, enabling them to handle rare or unseen codes, though they often produce unstructured outputs that complicate clinical grounding and calibration [24]. LLM embedding methods bridge this gap by using the representational power of LLMs while maintaining compatibility with traditional predictive modeling frameworks. By converting EHR data into structured text and embedding it with general-purpose LLMs, we enable the use of standard classifiers such as logistic regression to generate clinically meaningful predictions. This approach retains the calibration strengths of traditional models while benefiting from the rich contextual understanding of LLMs.
@@
This study has several limitations. First, our approach relies on a manually designed EHR serialization designed to capture medically relevant information, which may introduce bias when comparing against models trained on raw data. Additionally, LLM embedding model performance is sensitive to the specific content and instructions in the serialized text, which may limit reproducibility and generalization. While the models achieve strong predictive accuracy across tasks, including few-shot settings, their large parameter counts lead to longer computation time and higher resource demands. Moreover, the use of LLM embeddings requires training a separate downstream classifier, foregoing the native zero-shot or few-shot capabilities of LLMs. We also limited EHR serialization to 4,096 tokens due to runtime constraints, which may exclude relevant long-term historical data. Finally, evaluation was conducted on only two datasets, which may limit the generalizability of our findings across diverse healthcare systems.
@@
Future research should explore serialization-free approaches that allow LLMs to process raw EHR data directly, reducing biases introduced by manual text transformation. Integrating zero-shot and few-shot prompting into the LLM-based embedding framework could further enhance model flexibility and reduce dependency on downstream training. Extending the effective context window beyond 4,096 tokens will be important for capturing comprehensive patient histories. Additionally, developing techniques to distill large LLMs into smaller, more efficient models could improve their practical use in clinical settings. Expanding evaluations to real-world deployments and investigating how to combine the complementary strengths of domain-specific EHR models and general-purpose LLMs will be essential for building robust, scalable EHR foundation models.
@@ 4 Methods
4.1 EHRSHOT Database and Prediction Task
The EHR data used in our experiments is from the EHRSHOT benchmark for few-shot evaluation of EHR foundation models [31]. We obtained version 2.1 of the dataset, which is accessible via gated access under a research data use agreement. This dataset comprises longitudinal records for 6,739 patients, 921,499 visits, and 41,661,637 clinical events collected between 1990 and February 8, 2023. Each clinical event is linked to a specific patient and includes information such as start time, end time, a semantic code, a value, unit, visit ID, and the corresponding OMOP source table. We used the official ehrshot-benchmark repository111GitHub repository: https://github.com/som-shahlab/ehrshot-benchmark as a starting point to design our experiments, enabling us to build on existing functionalities and to facilitate comparisons with prior methods. The benchmark uses the Framework for Electronic Medical Records (FEMR) 222GitHub repository: https://github.com/som-shahlab/femr, which provides Python classes for efficient loading and processing of EHR data. All extensions and experiments conducted for this paper have been made publicly available via our GitHub repository: https://github.com/stefanhgm/ehrshot-benchmark. The EHRSHOT benchmark defines a rigorous evaluation including 15 clinical prediction tasks categorized into four groups: operational outcomes, anticipating lab test results, assignment of new diagnoses, and anticipating chest X-ray findings [31]. Task labels are derived from clinical events; hence, a single patient could contribute multiple labels per task, resulting in significant variations in task-specific sample sizes. For instance, frequent events like lab tests provide considerably more examples compared to rarer events such as new diagnoses. The benchmark focuses on analyzing model performance in a few-shot setting, which is particularly relevant for large pre-trained foundation models [14] due to their ability to generalize from limited training data. To this end, the benchmark defines evaluation settings with a constrained number of training and validation examples. Specifically, for k in 1, 2, 4, 8, 12, 16, 24, 32, 48, 64, 128, the benchmark uses k positive and k negative training examples, along with k positive and k negative validation examples, to train and tune supervised classifiers. Testing is always performed on the full set of examples. The classifiers evaluated within the EHRSHOT framework include logistic regression, random forests, and GBMs [41]. Performance is reported using the AUROC and the AUPRC. For few-shot settings, we average the results over five runs with different seeds and compute bootstrapped 95% confidence intervals [31]. Macro averages are reported for each task group, and an overall macro average is provided across all groups.
@@ 4.2 EHR Text Serialization
To use LLM embedding models for representing EHR records, we serialized the records into textual formats. We limited the serialization length to 4,096 tokens (approximately 16,000 characters) due to computational constraints (see Section 4.6). Two experimental runs were conducted with serializations extending up to 8,192 tokens. The primary goal was to create a detailed and informative serialization requiring minimal preprocessing while ensuring that medically relevant information appeared early in the text. This approach mitigated truncation risks in lengthy records, preserving critical details even when older entries were omitted. To convert the visits and clinical events in the EHRSHOT dataset into text, we used the semantic information embedded in the dataset. Each clinical event was labeled using the format “ontology/code”. EHRSHOT provided a set of prepared ontologies for resolving concept codes into their descriptions, which we incorporated. An analysis was performed to evaluate the use of these ontologies for all events of a subset of 200 patients across the task groups for operational outcomes and new diagnoses, covering 2,968 labels. We identified the following ontologies: Logical Observation Identifiers Names and Codes, SNOMED, RxNorm, CPT4, Domain, CARE_SITE, RxNorm Extension, Medicare Specialty, ICD10PCS, CMS Place of Service, Cancer Modifier, ICD9Proc, CVX, ICDO3, HCPCS, OMOP Extension, Condition Type. We excluded ontologies containing only a single value (Domain, Medicare Specialty, CMS Place of Service, OMOP Extension, and Condition Type). Codes of the ontologies CPT4, CARE_SITE, ICD10PCS, Cancer Modifier, CVX, and ICDO3 were not resolved with the provided ontologies. Cancer Modifier codes contained UICC cancer stages that we parsed manually. For CPT4, ICD10PCS, and CVX we used custom mapping files that we manually added.333We downloaded CPT4 from https://gist.github.com/lieldulev/439793dc3c5a6613b661c33d71fdd185, ICD10PCS from https://hcup-us.ahrq.gov/toolssoftware/procedureicd10/procedure_icd10_archive.jsp, and CVX from https://www2a.cdc.gov/vaccines/iis/iisstandards/vaccines.asp?rpt=cvx. We excluded CARE_SITE and ICDO3 since we found no way to resolve these to useful descriptions.
@@
Various approaches exist for serializing structured data, including row-by-row serialization [42], template-based methods [43], or structured data formats like JSON, HTML, and Markdown [44]. We used Markdown due to its minimal overhead and overall benefits of using a structured data input format for LLMs [45]. To harmonize dates, all timestamps were normalized relative to January 1, 2024, designated as the prediction reference time. Serialization began with patient demographics, typically the first event for each patient, where birthdates were converted into ages (in years) for simplicity. Since 65% of the dataset comprised time-series data encoded via LOINC, which we found imbalanced, we aggregated LOINC-coded events. Using the same patient subset as the ontology analysis, we identified the most frequent codes and categorized them into vital signs, body metrics, and lab values, selecting 24 key medical concepts. To avoid duplicates, we merged synonymous codes (see Table 7). The last three values of each concept are presented, and we filtered implausible values. To further enrich the text representation, we manually added default units and assessments (low, normal, high) based on standard ranges (see Table 7). Following the aggregated data, a summary of all visits was included to address the potential truncation of older visits. Events not associated with visits were then presented, using the same aggregation logic to display the last three values where applicable. Finally, a detailed chronological presentation of all visits was included, with events categorized into conditions (SNOMED, Visit, Cancer Modifier, CVX, HCPCS), medications (RxNorm, RxNorm Extension), and procedures (CPT4, ICD10PCS, ICD9Proc).